OpenStudy (anonymous):

for(x)=1/4x-5 find (fof-1) (-7)

OpenStudy (anonymous):

f (x) = 1/ 4x-5 f-1(x) = 1+5x / 4x f-1(-7)= 17/14 f(17/14)= -7 (fof-1) = -7